log(base 5)(X+3)-log(base 5)X=2
solve for x

Answer by ewatrrr(24785)

 
Hi
Familarize Yourself with the Properties:






5^2 = (x+3)/x
25x = x+3
24x = 3
x = 3/24 = 1/8